1)the length of a rectangular rug is 1 ft more than the width and a diagonal of the rug is 5 ft. Find the length and the width. 2) one leg of right triangle is 7 cm less than the length of the other leg. The length of the hypotenuse is 13 cm. Find the lengths of the legs
Answer by checkley77(12844)

1) L=W+1
L^2+W^2=5^2
(W+1)^2+W^2=25
W^2+2W+1+W^2=25
2W^2+2W+1-25=0
2W^2+2W-24=0
2(W^2+W-12)=0
(W+4)(W-3)=0
W-3=0
W=3 ANS.
L=3+1=4 ANS.
PROOF:
4^2+3^2=5^2
16+9=25
25=25